# Incremental Rotation Averaging Revisited and More: A New Rotation Averaging Benchmark

### v3

@@ Abstract
In order to further advance the accuracy and robustness of the incremental parameter estimation-based rotation averaging methods, in this paper, a new member of the Incremental Rotation Averaging (IRA) family is introduced, which is termed as IRAv4. As the most significant feature of the IRAv4, a task-specific connected dominating set is extracted to serve as a more reliable and accurate reference for rotation global alignment. In addition, to further address the limitations of the existing rotation averaging benchmark of relying on the slightly outdated Bundler camera calibration results as ground truths and focusing solely on rotation estimation accuracy, this paper presents a new COLMAP-based rotation averaging benchmark that incorporates a cross check between COLMAP and Bundler, and employ the accuracy of both rotation and downstream location estimation as evaluation metrics, which is desired to provide a more reliable and comprehensive evaluation tool for the rotation averaging research. Comprehensive comparisons between the proposed IRAv4 and other mainstream rotation averaging methods on this new benchmark demonstrate the effectiveness of our proposed approach.
Index Terms:
Large-scale rotation averaging, task-specific connected dominating set, rotation averaging benchmarking.I Introduction
@@
Image-based large-scale scene 3D reconstruction is a fundamental task in computer vision community, which has been widely investigated in recent years [1, 2, 3, 4]. As its core step, Structure from Motion (SfM) [5, 6, 7, 8, 9] aims to simultaneously recover camera pose and scene structure given pair-wise image feature matches. According to the camera pose initialization scheme, SfM methods could be roughly divided into incremental and global ones. The camera poses are sequentially estimated in incremental SfM via a iterative optimization pipeline [6, 8], but are simultaneously solved in global SfM by the motion averaging methodology [5, 9]. Compared with incremental SfM methods, global ones feature fewer optimization iterations and estimation parameters, which makes them with more advantage in theory and greater potential for application when dealing with increasingly larger reconstruction scene scales.
@@
Motion averaging, which takes relative camera motions (rotations and translations) as input and produces absolute camera poses (orientations and locations), is the primary technique used for camera pose recovery in global SfM. Due to scale ambiguity of the relative translation estimated via essential matrix decomposition [10], motion averaging is conducted in the manner of first rotation and then translation averaging in most cases. As the former of the above two phases, rotation averaging [11] directly influences the effect of the subsequent translation averaging phase, and even all the remaining 3D reconstruction procedure, e.g. multi-view triangulation for scene recovery and global Bundle Adjustment (BA) for parameter optimization. Though increasing attention has been drawn in recent years [12, 13, 14, 15, 16, 17, 18, 19, 20], the problem of rotation averaging is still far from being solved due to the large scale, imbalanced connectivity, and high-level noise in the Epipolar-geometry Graph (EG).
@@
Recently, a series of rotation averaging methods [21, 22, 23, 24] are proposed based on the ideology of incremental parameter estimation stemmed from incremental SfM. As the primitive and primary method, Incremental Rotation Averaging (IRA) [21] performs incremental absolute rotation computation and relative rotation outlier filtering simultaneously, by which the rotation estimation accuracy and robustness are both guaranteed. To further enhance the efficiency and scalablility, IRA++ [22] is proposed, where the input EG is clustered to construct several low-level intra-sub-EGs and a high-level inter-one. Then IRA is performed on all the (intra- and inter-) sub-EGs to achieve local estimation and global alignment of absolute rotations, respectively. In order to achieve a task-specific EG clustering for better rotation averaging performance, IRAv3 [23] is presented, where the cluster affiliation of each camera is dynamically determined with its absolute rotation (in the local coordinate system of intra-sub-EG) simultaneously estimated. To accomplish a better global alignment of the local rotation estimates, inspired by Jiang et al. [25], IRAv3+ [24] is introduced. Instead of performing a cluster-level rotation averaging, which is done in both IRA++ and IRAv3, multiple Connected Dominating Sets (CDSs) are randomly extracted to serve as the reference for rotation local-to-global alignment.
@@
However, in this paper we argue that the accuracy and robustness of the IRA series described above could be advanced one step further. Though the effectiveness of the cluster-based pipeline with task-independent CDS serving as global reference has been demonstrated in both IRAv3+ [24] and Jiang et al. [25], we believe that with a task-specific CDS extracted, more reliable global reference construction together with more accurate camera pose globalization would be further achieved. Based on the above analysis, a novel rotation averaging method termed as IRAv4 is proposed in this paper, which is built upon IRAv3+. The major difference between them lies in that instead of first extracting multiple CDSs and then estimating the absolute rotations in the CDSs’ local coordinate system by leveraging IRA, which is done in IRAv3+, the (task-specific) CDS is extracted by incrementally selecting the Next-Best Vertex (NBV) to maximize the supports from the currently extracted ones in the CDS, together with its absolute rotation (in the local coordinate system of CDS) simultaneously estimated.
Moreover, we find two issues that need to be addressed when evaluating the performance of most existing rotation averaging methods [26, 27, 28, 29, 16, 21, 30, 19, 31, 22, 23, 24]. Firstly, the evaluation is mostly performed based on the 1DSfM [32] dataset with the camera calibration results of the slightly outdated Bundler [33] serving as ground truth. And secondly, the evaluation mostly focuses on the camera rotation estimation accuracy only while ignoring the influence of the estimated rotations to the downstream task, i.e. translation averaging. To deal with these, a new COLMAP-based [6] rotation averaging benchmark is rebuilt upon the 1DSfM dataset in this paper, where the EG of each test data in the 1DSfM dataset111https://www.cs.cornell.edu/projects/1dsfm/ is regenerated by leveraging COLMAP222https://demuc.de/colmap/ and OpenCV333https://opencv.org/ libraries. To provide a more reliable ground-truth source, only the camera poses pass the cross check between COLMAP and Bundler are employed in this new benchmark. In order to additionally evaluate the effectiveness of the estimated rotations for the downstream translation averaging task, in our new benchmark, the rotations estimated by different rotation averaging methods, together with the relative translations in the regenerated EG, are feed into a well-established translation averaging method, BATA [34], to obtain absolute camera locations for location estimation accuracy evaluation. Based on the new benchmark, comprehensive evaluation is performed on the proposed IRAv4 and several currently mainstream rotation averaging methods, including the existing IRA series [21, 22, 23, 24], and some other methods [35, 26, 27, 30, 31, 28]. Among all these methods for performance evaluation, the proposed IRAv4 in this paper achieves state of the art in both rotation estimation and the downstream location estimation on the new benchmark, by which its effectiveness is demonstrated.
The main contributions of this paper are threefold:
1) A novel cluster-based rotation averaging pipeline is proposed, where a task-specific CDS is extracted to serve as the global reference for rotation local-to-global alignment. 2) A new rotation averaging benchmark based on the 1DSfM dataset and the COLMAP library is presented, where both the ground truth source and the performance evaluation metrics are redefined for a more reliable and comprehensive evaluation. 3) A comprehensive evaluation is conducted between the proposed IRAv4 and several mainstream rotation averaging methods on the new benchmark, where state-of-the-art performance of IRAv4 is achieved to demonstrate its effectiveness.
II Brief Description on the IRA Series
@@
III The Proposed IRAv4 Method
@@
III-A Task-Specific CDS-Based Global Reference Construction
Similar to Jiang et al. [25] and IRAv3+ [24], CDS-based sub-EG is also used to serve as a global reference for camera pose alignment in IRAv4. The CDS problem in graph theory asks for a minimum-size and connected subset of vertices with the following property which accounts for the concept of dominating: Each vertex is required to either be in the CDS, or adjacent to some vertex of the CDS. The CDS problem is usually solved based on an approximation algorithm [37], which roughly proceeds as follows:
@@ III-A Task-Specific CDS-Based Global Reference Construction
the above optimization problem and the other ones in the rest of this paper are solved by the Ceres Solver library444http://www.ceres-solver.org/, and the triplet with the largest selection reward defined in the following, together with the optimized rotations are served as the initial seed construction:
@@
and the edges(vertices) () and () for the vertex ’s rotation pre-computation and the iteration step’s NBV determination are selected by:
@@
where is the rotation initialization of the vertex . Then, the absolute rotation of is locally optimized by:
@@
III-B Other Key Steps Stemming from the IRAv3+
@@
On-the-fly EG clustering and local rotation averaging: This step is first adopted in IRAv3, which is similar in principle to the iteration step of the task-specific CDS-based global reference construction in IRAv4 (cf. Eq. 5 to Eq. 12). To achieve this, for each unaffiliated vertex, the edge set between it and the currently affiliated vertex set of each cluster is first obtained. Then, by leveraging these edge sets, the supporting edge set of each unaffiliated vertex within each cluster could be obtained by absolute rotation pre-computation (cf. Eq. 5) and relative rotation re-computation (cf. Eq. 6). On this basis, the next-best affiliated vertex, together with its affiliated cluster and initialized rotation are simultaneously determined by supporting set global maximization (cf. Eq. 7 and Eq. 8). Then, the rotation estimation are locally (only on the next-best affiliated vertex) or globally (on all the vertices of the affiliated cluster) optimized (cf. Eq. 9 to Eq. 12). The above procedure is itratively performed until all the vertices are affiliated, together with their rotations estimated.
@@
Common vertices-guided alignment rotation estimation: After reference construction and EG clustering, cluster-to-reference rotation is estimated for local absolute rotation alignment. As observed in IRAv3+ that the rotation averaging-based absolute rotation estimates are more accurate and reliable than those of the essential matrix decomposition-based relative ones, a common vertices-guided alignment rotation estimation method is proposed. For a particular cluster, either one common vertex or one shared edge between it and the reference could induce an estimate of the cluster-to-reference alignment rotation. As the common vertices-induced ones are with higher priority, they are served as guidance for alignment rotation estimation. Specifically, for each vertex-induced estimate, one could obtain its supporters from the edge-induced ones, and the one with most supporters is used as the initialization of the alignment rotation, which is further optimized by leveraging the constraints provided by its supporters. Readers may refer to the original paper of IRAv3+ [23] for more details.

### v4

@@ Abstract
In order to further advance the accuracy and robustness of the incremental parameter estimation-based rotation averaging methods, in this paper, a new member of the Incremental Rotation Averaging (IRA) family is introduced, which is termed as IRAv4. As its most significant feature, a task-specific connected dominating set is extracted in IRAv4 to serve as a more reliable and accurate reference for rotation local-to-global alignment. This alignment reference is incrementally constructed, together with the absolute rotations of the vertices belong to it simultaneously estimated. Comprehensive evaluations are performed on the 1DSfM dataset, by which the effectiveness of both the reference construction method and the entire rotation averaging pipeline proposed in this paper is demonstrated.
keywords:
[…]
1 Introduction
@@
Image-based large-scale scene 3D reconstruction is a fundamental task in computer vision community, which has been widely investigated in recent years [1, 2, 3, 4]. As its core step, Structure from Motion (SfM) [5, 6, 7, 8, 9, 10, 11] aims to simultaneously recover camera pose and scene structure given pair-wise image feature matches. According to the camera pose initialization scheme, SfM methods could be roughly divided into incremental and global ones. The camera poses are sequentially estimated in incremental SfM via a iterative optimization pipeline [6, 8], but are simultaneously solved in global SfM by the motion averaging methodology [5, 9]. Compared with incremental SfM methods, global ones feature fewer optimization iterations and estimation parameters, which makes them with more advantage in theory and greater potential for application when dealing with increasingly larger reconstruction scene scales.
@@
Motion averaging, which takes relative camera motions (rotations and translations) as input and produces absolute camera poses (orientations and locations), is the primary technique used for camera pose recovery in the global SfM. Due to scale ambiguity of the relative translation estimated via essential matrix decomposition [12], motion averaging is conducted in the manner of first rotation and then translation averaging in most cases [13, 14, 15, 16, 9, 17]. As the former of the above two phases, rotation averaging [18] directly influences the effect of the subsequent translation averaging phase [19], and even all the remaining 3D reconstruction procedure, e.g. multi-view triangulation for scene recovery and global Bundle Adjustment (BA) for parameter optimization. Though increasing attention has been drawn in recent years [20, 21, 22, 23, 24, 25, 26, 27, 28], the problem of rotation averaging is still far from being solved due to the issues of large scale, imbalanced connectivity, and high-level noise of the Epipolar-geometry Graph (EG).
@@
Recently, a series of rotation averaging methods [29, 30, 31, 32] are proposed based on the ideology of incremental parameter estimation stemmed from the incremental SfM. As the primitive and primary method, Incremental Rotation Averaging (IRA) [29] performs incremental absolute rotation computation and relative rotation outlier filtering simultaneously, by which the rotation estimation accuracy and robustness are both guaranteed. To further enhance the efficiency and scalablility, IRA++ [30] is proposed, where the input EG is clustered to construct several low-level intra-sub-EGs and a high-level inter-one. Then IRA is performed on all the (intra- and inter-) sub-EGs to achieve local estimation and global alignment of absolute rotations, respectively. In order to achieve a task-specific EG clustering for better rotation averaging performance, IRAv3 [31] is presented, where the cluster affiliation of each camera is dynamically determined with its absolute rotation (in the local coordinate system of intra-sub-EG) simultaneously estimated. To accomplish a better global alignment of the local rotation estimates, inspired by the method of WCDS [33] proposed by Jiang et al., IRAv3+ [32] is introduced. Instead of performing a cluster-level rotation averaging, which is done in both IRA++ and IRAv3, multiple Connected Dominating Sets (CDSs) are randomly extracted to serve as the reference for rotation local-to-global alignment.
@@
However, in this paper we further argue that the accuracy and robustness of the IRA series described above could be further advanced. Though the effectiveness of the cluster-based pipeline with task-independent CDS serving as global reference has been demonstrated in both IRAv3+ [32] and WCDS [33], we believe that with a task-specific CDS extracted, more reliable global reference construction together with more accurate camera pose globalization would be further achieved. Based on the above analysis, a novel rotation averaging method termed as IRAv4 is proposed in this paper, which is built upon IRAv3+. The major difference between them lies in that instead of first extracting multiple CDSs and then estimating the absolute rotations in the CDSs’ local coordinate system by leveraging IRA, which is done in IRAv3+, the (task-specific) CDS is extracted by incrementally selecting the Next-Best Vertex (NBV) to maximize the supports from the currently extracted ones in the CDS, together with its absolute rotation (in the local coordinate system of CDS) simultaneously estimated. Our IRAv4 is evaluated on the 1DSfM dataset and compared with several other mainstream rotation averaging methods to demonstrate its effectiveness and advantages.
2 Related works
Except for the IRA series, there are mainly three kinds of large-scale rotation averaging methods, including the robust loss-based [34, 35, 36], the outlier filtering-based [37, 38, 39], and the deep learning-based [40, 41, 42] ones, which are briefly reviewed in the following.
2.1 Robust loss-based rotation averaging
The robust loss-based methods [34, 35, 36] seek to design robust loss functions to achieve a robust estimation of the absolute camera rotations in the presence of relative rotation measurement outliers. Chatterjee et al. [34] developed a two-stage rotation averaging method. They firstly use the solution as an initialization to make the method robust to outliers. Then, an Iteratively Reweighted Least Squares (IRLS) approach based on the loss function is followed to achieve an efficient and accurate absolute rotation estimation. Shi et al. [35] proposed a novel robust optimization framework, Message Passing Least Squares (MPLS), as an alternative paradigm to IRLS for rotation averaging, and the advantages of MPLS over IRLS are theoretically analyzed. After that, based on the cyclic consistency constraint, Shi et al. [36] presented a novel quadratic programming framework, Detection and Estimation of Structural Consistency (DESC), which is less dependent to initialization and more simple in optimization. Though the above methods are robust to the relative rotation outliers to some extent, the optimization procedures in them still heavily depend on the initialization and are usually inefficient because of the relatively complicated problem formulations.
2.2 Outlier filtering-based rotation averaging
Compared with the robust loss-based methods, the outlier filtering-based ones [37, 38, 39], are more theoretically straightforward. Cui et al. [37] proposed a multiple Orthogonal Maximum Spanning Trees (OMSTs)-based rotation averaging method. Multiple OMSTs are selected from the input EG to obtain a set of densely connected and more accurate relative rotation measurements for robust rotation averaging [34]. Gao et al. [38] presented a Hierarchical RANSAC-based Rotation Averaging (HRRA) method. The absolute rotations are estimated based on the thought of RANSAC with random spanning tree of the input EG serving as the minimal set for model estimation. In addition, graph clustering is employed to downsize the spanning trees as the RANSAC-based parameter estimation is sensitive to the size of the minimal set. Lee et al. [39] introduced the algorithm of HARA, which incrementally adds constraints based on the size of the supporting camera triplets to construct EG’s spanning tree and remove outliers in relative rotation measurements, thus achieving a hierarchical and robust initialization of camera absolute rotation. Though effective outlier filtering and accurate rotation estimation could be achieved in most cases, we find that the performance of the above methods exhibit some instability when dealing with the EGs with relatively high noise level.
2.3 Deep learning-based rotation averaging
Recently, several deep learning-based methods have been proposed dealing with the rotation averaging problem by leveraging the techniques of graph neural networks [40, 41, 42]. Purkait et al. [40] proposed NeuRoRA, a two-stage graph neural networks for rotation averaging. The first stage is a view-graph cleaning network to detect relative rotation measurement outliers. The second stage is a fine-tuning network to refine the absolute orientations. Yang et al. [41] presented Multi-Source Propagator (MSP), which takes appearance information as input and introduces a differentiable Shortest Path Tree (SPT) method to achieve a robust initialization result. The initialization is further improved through non-learnable iterative edge reweighting. Li et al. [42] proposed the method of Rotation Averaging Graph Optimizer (RAGO), a learnable iterative graph optimizer minimizing a gauge-invariant cost function with an edge rectification strategy to mitigate the effect of inaccurate measurements. However, these deep learning-based methods depend heavily on the training data, which makes their generalization and scalability poor in novel situations.
3 Brief description on the previous IRA series
@@
4 The proposed IRAv4 method
@@
4.1 Task-specific alignment reference construction
Similar to WCDS [33] and IRAv3+ [32], CDS-based sub-EG is also used to serve as a global reference for camera pose alignment in IRAv4. The CDS problem in graph theory asks for a minimum-size and connected subset of vertices with the following property which accounts for the concept of dominating: Each vertex is required to either be in the CDS, or adjacent to some vertex of the CDS. The CDS problem is usually solved based on an approximation algorithm [44], which roughly proceeds as follows:
@@ 4.1 Task-specific alignment reference construction
the above optimization problem and the other ones in the rest of this paper are solved by the Ceres Solver library, and the triplet with the largest selection reward defined in the following, together with the optimized rotations are served as the initial seed construction:
@@
and the edges(vertices) () and () for the pre-computation of vertex ’s rotation and the NBV determination of iteration step are selected by:
@@
where is the initialization of the vertex . Then, the absolute rotation of is locally optimized by:
@@
4.2 Other key steps stemming from the IRAv3+
@@
On-the-fly EG clustering and local rotation averaging: This step is first adopted in IRAv3, which is similar in principle to the iteration step of the task-specific CDS-based global reference construction in IRAv4 (cf. Eq. 5 to Eq. 12). To achieve this, for each unaffiliated vertex, the edge set between it and the currently affiliated vertex set of each cluster is first obtained. Then, by leveraging these edge sets, the supporting edge set of each unaffiliated vertex within each cluster could be obtained by absolute rotation pre-computation (cf. Eq. 5) and relative rotation re-computation (cf. Eq. 6). On this basis, the next-best affiliated vertex, together with its affiliated cluster and initialized rotation are simultaneously determined by supporting set global maximization (cf. Eq. 7 and Eq. 8). Then, the rotation estimation are locally (only on the next-best affiliated vertex) or globally (on all the vertices of the affiliated cluster) optimized (cf. Eq. 9 to Eq. 12). The above procedure is iteratively performed until all the vertices are affiliated, together with their rotations estimated.
@@
Common vertices-guided alignment rotation estimation: After reference construction and EG clustering, cluster-to-reference rotation is estimated for local rotation alignment. As observed in IRAv3+ that the rotation averaging-based absolute rotation estimates tend to be more accurate and reliable than those of the essential matrix decomposition-based relative ones, a common vertices-guided alignment rotation estimation method is proposed. For a particular cluster, either one common vertex or one shared edge between it and the reference could induce an estimate of the cluster-to-reference alignment rotation. As the common vertices-induced ones are with higher priority, they are served as guidance for alignment rotation estimation. Specifically, for each vertex-induced estimate, one could obtain its supporters from the edge-induced ones, and the one with most supporters is used as the initialization of the alignment rotation, which is further optimized by using the constraints provided by its supporters. Readers may refer to the paper of IRAv3+ [31] for more details.
[…]
Data &
[…]
ALM
ELS
GDM
MDR
MND
NYC
PDP
PIC
ROF
TOL
TFG
USQ
VNC
YKM
[…]